500 g of water and 100g of ice at 0 degree are in a calorimeter whose water equivalent is 40 g 10g of steam at 100 degree is added to it then water in the calorimeter is (latent heat of ice = 80 cal/g latent heat of steam = 540 cal/g)

Grade:12th Pass

1 Answers

Arun
3 years ago
Here, the energy released by the condensation of vapour is not sufficient to melt the whole ice. So, the amount of energy released by the water vapour, when it condensed and comes to 00C is,
 
E=(mvapour)(Lvapour)+(mvapour)(Svapour)×100⇒E=10×540+10×1×100⇒E=6400 cal
 
So, let m mass of the ice is m'. Therefore,
 
m'Lice=6400 ⇒m'=640080⇒m'=80 gm
Therefore, the total amount of water in the capillary tube is,
 
m0=500+80+10 ⇒m0=590 gm